OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11

sam (.sam.):

3x − 6y = 3 7x − 5y = −11 ----------------------- 3x = 3+ 6y x = (3+ 6y)/3 -------------------- 7x − 5y = −11 7 (3+ 6y)/3 − 5y = −11 (21)(1+2y)/3-5y=-11 7(1+2y)-5y=-11 7+14y-5y=-11 7+9y=-11 9y=-18 y=-2 ------------------ x = (3+ 6y)/3 x=(3+6(-2))/(3) x=(3-12)/(3) x=(-9)/(3) x=-3

OpenStudy (swag):

X=-3